40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 311     Entry time: Thu Jan 10 20:42:54 2019
Author: Koji 
Type: Optics 
Category: Characterization 
Subject: FSR / HOM Test of OMC SN002 

OMC SN002 = Former LHO OMC which CM1 was destroyed by the lock loss pulse in 2016. This OMC needs to be optically tested before storage.

The test items:

  • [done] FSR measurement with offset PDH locking (FM->AM conversion)
  • [done] FSR/finesse measurement with the EOM RFAM injection
  • [done] TMS measurement with input miaslignment and the trans RFPD misalignment: with no PZT offset
  • [done] TMS measurement with input miaslignment and the trans RFPD misalignment: with PZT offsets
     
  • PZT response
  • Mirror cleaning
  • Power budget
  • Diode alignment: shim height
  • PD/QPD alignment
ELOG V3.1.3-